A commercial method of manufacturing hydrogen involves the reaction of iron and
steam. (a) How many grams of H2 can be produced from 42.7 g Fe and an excess of
H20(g) (steam)? (b) How many grams of H2O are consumed in the conversion of
63.5 g Fe to Fe3O4
1
Expert's answer
2020-12-21T03:47:12-0500

a)Fe+H2O=FeO+H2amountofsubstanceofFen=m/Ar=42.7/56=0.76moln(H2)=n(Fe)V(H2)=nVm=0.76∗22.4=17.024l.b)3Fe+4H2O=4H2+Fe3O4.n(Fe)=m/Mr=63.5/56=1.13mol.n(H2O)=4/3n(Fe)=4/3∗1.13=1.5mol.m(H2O)=nMr=1.5∗18=27g